OpenStudy (anonymous):

Rita bought some large packets of salt weighing 500 grams each and small packets of salt weighing 200 grams each. If she has 8 packets altogether and they weigh 2200 grams in total, how many of each kind does she have?

OpenStudy (anonymous):

L + S = 8 500L + 200S = 2200 L = 8 - S 500(8-S) + 200S = 2200 4000 - 500S + 200S = 2200 300S =1800 S = 6 L = 8 - 6 = 2